# A machine-learning classifier for the postmerger remnant of binary neutron stars

### v1

@@ III.1.1 Real events: GW170817 and GW190425
In the following, we employ our classifier to predict the merger outcome of GW170817 and GW190425 based on the samples values for the parameters as inferred from the analysis of the measured GW signal. All the predictions regarding these two real events, including results for the other classifiers discussed in the upcoming sections, are summarized in Table 2.
@@
with $n_{i}$ being the number of samples predicting the scenario $i$. We find that the probability of prompt collapse to BH (PCBH) and formation of a NS remannt (RNS) are $p_{\rm\textsc{pcbh}}\sim 40\%$ and $p_{\rm\textsc{rns}}\sim 60\%$, respectively.
However, this result is derived from the GW data alone, and more stringent constraints are obtained if we include additional information about the EoS. In Ref. [95], the GW170817 and GW190425 data are re-analyzed with a prior on the EoS that takes into account, among others, constraints derived from chiral effective field theory, perturbative quantum chromodynamics, and the minimum values for $M_{\rm\textsc{tov}}$ inferred from heavy radio pulsars. If we consider the posteriors for GW170817 produced in this work with such additional constraints (GW170817+EoS), Classifier A predicts a much larger probability for the formation of a remnant, with $p_{\rm\textsc{pcbh}}\sim 9\%$ and $p_{\rm\textsc{rns}}\sim 91\%$. The results obtained from the two sets of samples are consistent with the findings in Ref. [39], where, with two different methods, the probability of prompt collapse to BH for GW170817 is predicted to be $\sim 50-70\%$, but below $\sim 10\%$ when mass information from the heavy pulsars is included.
@@
Additionally, Ref. [95] re-analyzed GW170817 incorporating also information from the EM counterpart [97], the kilonova AT2017gfo [98] (GW170817+EoS+KN) and the short gamma-ray burst GRB170817A [99] (GW170817+EoS+KN+GRB). For the posteriors produced including EM data, the predicted probability of formation of a NS remnant increases further, up to $99.9\%$ (see Table 2).
Figure 5: Distribution of $M_{\rm tot}$ and $\tilde{\Lambda}$ samples values produced in Ref. [95] for GW170817+EoS, where the different colors indicate that Classifier A predicted for that sample a prompt collapse to BH (teal) or formation of a NS remnant (orange). Note that we show only the $\tilde{\Lambda}$ and $M_{\rm tot}$ values of the samples since they are the most relevant ones for the classifier, but the input features include also the sample $q$ and $\chi_{\rm eff}$ values, which are not shown here for simplicity.
@@ III.1.2 Prompt-collapse threshold and effect of parameters
Because of the impact on possible electromagnetic counterparts, it is particularly important to predict whether a BNS system underwent prompt collapse or not. As mentioned in Sec. I, if the mass of the remnant exceeds the maximum mass supported with differential rotation, the system is bound to a prompt collapse. This mass threshold value $M_{\rm thr}$ is therefore considered the key parameter to determine whether the system forms a NS remnant or not. However, $M_{\rm thr}$ depends on the EoS, thus finding a connection between the remnant mass and the probability of prompt collapse is not straightforward. Refs. [35, 20] noticed that $M_{\rm thr}$ can be expressed as a function of $M_{\rm\textsc{tov}}$ as $M_{\rm thr}=\kappa\cdot M_{\rm\textsc{tov}}$, where $\kappa$ depends on the EoS. Subsequent studies improved this relation, taking into account also the effects of mass ratio and spins [36, 28, 39, 37, 79, 38, 21].
@@
In order to understand the effect of mass ratio, we generate a grid of points with $M_{\rm tot}\in[2.4,3.5]\,M_{\odot}$, $q\in[1.,2.]$, and $\tilde{\Lambda}\in[10,2000]$. The central panel of Fig. 6 shows the distribution of prompt collapse and NS remnant predictions, assuming $\chi_{\rm eff}=0.0$. Besides the strong dependence on $\tilde{\Lambda}$, one can distinguish also the influence on $M_{\rm tot}$ and $q$, with higher values of both leading more likely to prompt collapse. While it is expected that larger values of total mass lead to a prompt collapse, understanding the effects of mass ratio is not straightforward. However, different studies showed that mass asymmetry decreases $M_{\rm thr}$ for most EoSs [^5] Ref. [79] pointed out that, for some specific EoSs, intermediate values of $q$ can lead to the opposite effect. (see, e.g., Refs. [28, 38, 36, 37]), consistent with our model’s behavior.
@@ III.2 Classifier B
Next, we build a classifier to distinguish between three different scenarios: prompt collapse to BH (with probability dubbed again $p_{\rm\textsc{pcbh}}$), formation of a HMNS that collapses within the simulation (with probability $p_{\rm\textsc{hmns}}$), or formation of a remnant that does not collapse within the simulation time ($p_{\rm\textsc{nc}}$). As explained in Sec. II.1, our dataset includes simulations whose remnant time, $\tau_{\rm BH}$, is registered as “greater” than the remaining simulation time after the merger, $t_{\rm sim}$. However, for some of them, this remaining time is very short, of the order of a few milliseconds. These cases are not informative, since a configuration for which we only know that the collapse time is larger than 5 ms, for example, could both be a HMNS or belong to class (IV). Therefore, for Classifier B, as well as for Classifier C (see next Sec. III.3), we remove from the dataset those points with $\tau_{\rm BH}>t_{\rm sim}$, if $t_{\rm sim}<25$ ms. This new dataset comprises 318 configurations.
As for Classifier A, the dataset is split in 90% for training and 10% for validation (see Sec. II.2), and we optimize the model’s hyperparameters. The confusion matrices computed on the validation and complete set are shown in Fig. 7, while the classifier’s accuracy and $MCC$ are reported in Table 1. Although both $\alpha$ and $MCC$ are slightly lower than the values found for Classifier A, as expected given the increased complexity of the task, they are still large, above $90\%$.
@@ III.3 Classifier C
Finally, we create a classifier to distinguish between all the four classes, employing the modified dataset as for Classifier B. We follow the same procedure, first creating the training and validation set as explained in Sec. II.2, and then optimizing the model hyperparameters. From Fig. 8, showing the confusion matrices obtained on the validation set and the complete dataset, and Tab. 1, we see that the accuracy and $MCC$ are lower than the values found for the other two classifiers, as expected, but we still recover only a few misclassifications, with a validation accuracy $\sim 88\%$. In particular, we do not observe cases of “extreme” misclassification, i.e., in which a prompt collapse is classified as class (IV) or vice versa.
@@
IV Uncertainty on models predictions
In this section, we discuss how to quantify the uncertainty on our classifiers predictions.
For each input point $\mathbf{x}$, the classifier computes the probabilities for $\mathbf{x}$ to belong to each class $\omega_{k}$, $P(\omega_{k}|\mathbf{x})$, and its final prediction is that $\mathbf{x}$ belongs to the class with the largest probability. The likelihood of the preferred class provides the confidence, and can be used as a measure of the uncertainty of the prediction [101]. Figure 9 shows the distribution of the prediction confidence for Classifier A over the GW170817 samples. Regarding as an overall uncertainty estimate the mean of this distribution, and considering, for example, the results of Sec. III.1, we find that our model predicts GW170817 to have $60\%$ probability of formation of a NS remnant and $40\%$ prompt collapse with confidence $>99\%$. For the GW170817 samples obtained with the additional EoS and EM information (GW170817, GW170817+EoS, GW170817+EoS+KN, GW170817+EoS+KN+GRB), we obtain a prediction confidence $\sim 99\%$ for Classifier A, $\sim 94\%$ for Classifier B, and $83-90\%$ for Classifier C. This estimate of uncertainty basically corresponds to quantifying “how sure” the classifier is about its prediction, therefore if two or more classes have very similar probabilities the uncertainty will be larger. This explains why Classifier B and Classifier C yield larger prediction uncertainties, especially in the latter case, where the classifier predicts basically the same probability for case (II) and case (III) for GW170817+EoS+KN+GRB (see Table 2).
@@ IV Uncertainty on models predictions
It is important to highlight that the largest source of uncertainty comes from the NR simulations employed to train the models. Different physics (e.g., neutrinos treatments), settings, and resolutions can produce different $\tau_{\rm BH}$ even if the initial system configuration is the same.
In the event of a detection, we need to know how much we can trust the prediction of these classifiers, especially if it informs EM followups. In this case, the input $\mathbf{x}$ would consist of only one single point with the estimated parameters values provided by low-latency analyses, and not the full posterior obtained though the subsequent parameter estimation process. Therefore, apart from estimating the confidence of the prediction as explained above, we can compute the probability of that point being misclassified to further decide whether to trust the prediction or not. We can calculate the uncertainty of the prediction for a point $\mathbf{x}$ based on the prediction entropy, and compare this value with a threshold found for the specific classifier employed: if the prediction uncertainty for $\mathbf{x}$ is smaller than the threshold, we trust the classification; if, on the other hand, the uncertainty is greater than the threshold, the probability of the prediction being a misclassification is non-negligible, and therefore more investigations about the event are necessary. In particular, for Classifier A, B, and C, we find threshold values $T_{\rm\textsc{a}}\sim 0.3$, $T_{\rm\textsc{b}}\sim 0.06$, and $T_{\rm\textsc{c}}\sim 0.09$, respectively. The details about the uncertainty and the thresholds calculations are explained in Appendix A.
@@ V Conclusions
We built a classifier to predict the outcome of a BNS merger based on parameters that can be measured from the inspiral GW signal, $M_{\rm tot}$, $q$, $\tilde{\Lambda}$, and $\chi_{\rm eff}$. We employed the Gradient Boosted Decision Tree algorithm in sklearn and a dataset comprising publicly available data from NR simulations. We construct and train three different models: Classifier A to distinguish between prompt collapse to BH and formation of a NS remnant; Classifier B to discriminate between prompt collapse, formation of a HMNS that collapses during the simulation, and formation of a NS remnant that does not collapse within the simulation time; Classifier C, which further discerns between short-lived ($2\,{\rm ms}<\tau_{\rm BH}<5\,{\rm ms}$) and long-lived ($\tau_{\rm BH}>5\,{\rm ms}$) HMNSs. All of them yield very high values for both accuracy and the Matthews correlation coefficient (see Table 1), where the latter takes into account the distribution of correct classifications among the different classes. The lowest values obtained are $\alpha=88.2\%$ and $MCC=0.831$ for Classifier C, as expected given its more complex task.
@@
We investigate the outcome of GW170817 and GW190425 with the three classifiers, supplying as input the samples for $M_{\rm tot}$, $q$, $\tilde{\Lambda}$, and $\chi_{\rm eff}$ produced by parameter estimation analysis on the GW data. We conclude that GW170817 had a $\sim 60\%$ probability of forming a NS remnant, most likely a HMNS that collapsed in a few ms, while GW190425 collapsed promptly to a BH, with a $60-70\%$ probability. Moreover, we also applied our classifiers to the posteriors samples obtained in Ref. [95] including in the inference information about the EoS and, in the case of GW170817, the EM counterpart. In this case, the probability for GW170817 to have formed a HMNS becomes larger than $99\%$, with roughly the same probability to be short- and long-lived, while GW190425 has more than $98\%$ probability to have undergone prompt collapse (see Table 2 for details).
@@
For all the three classifiers, the most significant feature in determining the prediction is $\tilde{\Lambda}$, as expected since it is the parameter encoding the information about the EoS. Although the importance of $\chi_{\rm eff}$ is overall quite low, for classifier A we find that, in general, binaries with $\chi_{\rm eff}<0$ ($\chi_{\rm eff}>0$) are more likely to produce a prompt collapse to BH (NS remnant) with respect to the irrotational case (see Fig. 6). However, only few of the simulations included in our dataset correspond to spinning configurations, and we do not consider misaligned spins, thus more data are needed to systematically investigate the spin influence.
@@
For Classifier A, which discriminates between prompt collapse and formation of a NS remnant, in the case of the GW170817 event we observe a clear threshold value for $\tilde{\Lambda}$ between the two scenarios (see Fig. 5). Specifically, for the samples median values of $M_{\rm tot}$, $q$, and $\chi_{\rm eff}$, we find that prompt collapse occurs for $\tilde{\Lambda}<\tilde{\Lambda}_{\rm th}\sim 310$.
@@
Finally, we discuss the uncertainty on the predictions of our models based on the probability assigned to each class, showing that our results on the real events data yield a prediction confidence of roughly $90\%$ or higher.
@@
Knowing the remnant produced by a BNS merger could provide us with information about the EoS of NSs’ supranuclear-dense matter, the physical processes involved in the postmerger phase, and possible EM counterparts. Moreover, predicting whether a BNS system undergoes a prompt collapse to BH or produces a NS remnant is crucial to build postmerger models and to inform EM followup campaigns, especially for 3G detectors, for which a large number of BNS detections are expected. The classifiers developed in this work can be employed with any BNS detection, without the need of observing a postmerger signal. In fact, due to the high frequencies involved, the postmerger part of the GW signal is difficult to detect, and we expect to start observing it only with 3G detectors, but also in that case it could be weak [26, 29]. On the other hand, our models enable us to predict the merger outcome also for potential detections with Advanced Virgo and Advanced LIGO; moreover, with future generation detectors, the more precise measurements of the binary parameters from the inspiral signal will allow us to get even more precise predictions, even without including additional information.

### v2

@@ III.1.1 Real events: GW170817 and GW190425
In the following, we employ our classifier to predict the merger outcome of GW170817 and GW190425, based on the sample values for the parameters, as inferred from the analysis of the measured GW signal. All the predictions regarding these two real events, including results for the other classifiers discussed in the upcoming sections, are summarized in Table 3.
@@
with $n_{i}$ being the number of samples predicting the scenario $i$. We find that the probability of prompt collapse to BH (PCBH) and formation of a NS remnant (RNS) are $p_{\rm\textsc{pcbh}}\sim 40\%$ and $p_{\rm\textsc{rns}}\sim 60\%$, respectively.
However, this result is derived from the GW data alone, and more stringent constraints are obtained if we include additional information about the EoS. In Ref. [98], the GW170817 and GW190425 data are reanalyzed with a prior on the EoS that takes into account, among others, constraints derived from chiral effective field theory, perturbative quantum chromodynamics, and the minimum values for $M_{\rm\textsc{tov}}$ inferred from heavy radio pulsars. If we consider the posteriors for GW170817 produced in this work with such additional constraints (GW170817+EoS), Classifier A predicts a much larger probability for the formation of a remnant, with $p_{\rm\textsc{pcbh}}\sim 9\%$ and $p_{\rm\textsc{rns}}\sim 91\%$. The results obtained from the two sets of samples are consistent with the findings in Ref. [43], where, with two different methods, the probability of prompt collapse to BH for GW170817 is predicted to be $\sim 50-70\%$, but below $\sim 10\%$ when mass information from the heavy pulsars is included.
@@
Additionally, Ref. [98] reanalyzed GW170817 incorporating also information from the EM counterpart [100], the kilonova AT2017gfo [101] (GW170817+EoS+KN) and the short gamma-ray burst GRB170817A [102] (GW170817+EoS+KN+GRB). For the posteriors produced including EM data, the predicted probability of formation of a NS remnant increases further, up to $99.9\%$ (see Table 3).
Incorporating additional information in the analysis of GW170817 and GW190425 significantly improves the prediction accuracy mostly because the posterior probability densities obtained from GW data only are quite wide, especially for $\tilde{\Lambda}$. This translates into the presence of samples with combinations of masses and $\tilde{\Lambda}$ that are not compatible with most physical EoSs (for example, large values of $\tilde{\Lambda}$ for the quite large mass values in GW190425). Consequently, including information about the EoS places tighter constraints on the $\tilde{\Lambda}$ posteriors, producing more precise predictions. For future-generation detectors, we expect much tighter constraints on the parameters from the GW data alone (see, e.g., Ref. [103]), and therefore precise predictions also without additional information.
The inclusion of data from the kilonova and short gamma-ray burst considers information from the postmerger phase, but (i) it does not require the detection of a postmerger GW signal, and (ii) it shows an additional improvement of the analysis, it is not a requirement for the classifier usage. A potential independent detection of the postmerger EM counterpart could provide an interesting comparison with our classifiers’ prediction from GW data. For example, if the classifier predicts the formation of a remnant, but the EM measurements are indicative of a prompt collapse, this might point to a phase transition happening with the increased density in the postmerger phase. Nevertheless, to investigate this kind of effect, an extended study of all possible systematics would be needed, which requires a larger number of NR simulations that are not available at the moment.
Figure 5: Distribution of $M_{\rm tot}$ and $\tilde{\Lambda}$ sample values produced in Ref. [98] for GW170817+EoS, where the different colors indicate that Classifier A predicted for that sample a prompt collapse to BH (teal) or formation of a NS remnant (orange). Note that we show only the $\tilde{\Lambda}$ and $M_{\rm tot}$ values of the samples since they are the most relevant ones for the classifier, but the input features include also the sample $q$ and $\chi_{\rm eff}$ values, which are not shown here for simplicity.
@@ III.1.2 Prompt-collapse threshold and effect of parameters
Because of the impact on possible electromagnetic counterparts, it is particularly important to predict whether a BNS system underwent prompt collapse or not. As mentioned in Sec. I, if the mass of the remnant exceeds the maximum mass supported with differential rotation, the system is bound to endure a prompt collapse. This mass threshold value $M_{\rm thr}$ is therefore considered the key parameter to determine whether the system forms a NS remnant or not. However, $M_{\rm thr}$ depends on the EoS, thus finding a connection between the remnant mass and the probability of prompt collapse is not straightforward. Refs. [39, 24] noticed that $M_{\rm thr}$ can be expressed as a function of $M_{\rm\textsc{tov}}$ as $M_{\rm thr}=\kappa\cdot M_{\rm\textsc{tov}}$, where $\kappa$ depends on the EoS. Subsequent studies improved this relation, taking into account also the effects of mass ratio and spins [40, 32, 43, 41, 104, 42, 25].
@@
In order to understand the effect of mass ratio, we generate a grid of points with $M_{\rm tot}\in[2.4,3.5]\,M_{\odot}$, $q\in[1.,2.]$, and $\tilde{\Lambda}\in[85,2000]$. Although some of the points in this grid, for larger $M_{\rm tot}$ and lower $\tilde{\Lambda}$, are slightly outside the training-set range (see Tab. 1), here we intend to show a general qualitative behavior, and we do not expect the overall trends, i.e., larger values of $M_{\rm tot}$ and smaller values of $\tilde{\Lambda}$ leading more likely to a prompt collapse, to change over these slightly wider ranges.
The central panel of Fig. 6 shows the distribution of prompt collapse and NS remnant predictions, assuming $\chi_{\rm eff}=0.0$. Besides the strong dependence on $\tilde{\Lambda}$, one can also distinguish the influence on $M_{\rm tot}$ and $q$, with higher values of both leading more likely to prompt collapse. While it is expected that larger values of total mass lead to a prompt collapse, understanding the effects of mass ratio is not straightforward. However, different studies showed that mass asymmetry decreases $M_{\rm thr}$ for most EoSs [^5] Ref. [104] pointed out that, for some specific EoSs, intermediate values of $q$ can lead to the opposite effect. (see, e.g., Refs. [32, 42, 40, 41]), consistent with our model’s behavior.
@@
III.1.3 Application to low-latency data
As mentioned, one of the main applications of our classifier is to inform potential electromagnetic follow-up campaigns, especially with next-generation detectors, when we expect to observe a large number of BNS merger events. Currently, low-latency analyses provide point estimates of parameters obtained through template-based, matched-filtering searches. These estimates include the values of masses and spins, but not tidal deformabilities [^6] At present, the templates employed in BNS searches do not incorporate tidal deformability because it would imply a large additional computational cost, while, given its relatively small effect on the overall phase, it is not expected to significantly improve the search results, and therefore binary black hole waveforms are deemed sufficient.. Although this might change for future-generation detectors, since their increased sensitivity will require more and more accurate templates, for the moment we can marginalize over EoSs to estimate the value of $\tilde{\Lambda}$ from low-latency data (recently, EoS marginalization has been employed also for the prediction of kilonova curves in Ref. [105]). As an example, let us investigate what we would have obtained with the real events GW170817 and GW190425. For the point estimates of masses and spins released by the LVK public alerts (as reported in Table 1 in Ref. [106]), we consider the set of $10^{5}$ EoSs analyzed in Ref. [98], and for each one of them we compute the $\tilde{\Lambda}$ value corresponding to the low-latency mass estimates. At this point, we can either (i) take the $\tilde{\Lambda}$ estimate as their weighted mean, where the weights correspond to each EoS posterior probability as obtained through the analysis in Ref. [98] [^7] The posterior probabilities are provided by the web-interface in Ref. [107]. In particular, we chose a conservative estimate that included only information from the Black Widow J0952-0607, NICERJ0030+0451, NICERJ0740+6620, and GW170817 measurements; see Ref. [98] for details.; (ii) since the classifiers predict the probability for each class, and then choose as the output label the one with the highest probability, we can directly weigh the classes’ probabilities. With both these methods, for GW170817 Classifier A predicts the formation of a NS remnant, while for GW190425 a prompt collapse to BH.
Both these approaches provide us with a prediction in a few seconds. Moreover, a lot of effort is ongoing to develop methods to perform fast parameter estimation analyses (see, e.g., Refs. [108, 109, 110]), which are expected to deliver posterior samples within minutes, and hence, in the future, can be employed in the low-latency online analyses.
@@ III.2 Classifier B
Next, we build a classifier to distinguish between three different scenarios: prompt collapse to BH (with probability again dubbed $p_{\rm\textsc{pcbh}}$), formation of a HMNS that collapses within the simulation (with probability $p_{\rm\textsc{hmns}}$), or formation of a remnant that does not collapse within the simulation time ($p_{\rm\textsc{nc}}$). As explained in Sec. II.1, our dataset includes simulations whose remnant time, $\tau_{\rm BH}$, is registered as “greater” than the remaining simulation time after the merger, $t_{\rm sim}$. However, for some of them, this remaining time is very short, of the order of a few milliseconds. These cases are not informative, since a configuration for which we only know that the collapse time is larger than 5 ms, for example, could both be a HMNS or belong to class IV. Therefore, for Classifier B, as well as for Classifier C (see next Sec. III.3), we remove from the dataset those points with $\tau_{\rm BH}>t_{\rm sim}$, if $t_{\rm sim}<25$ ms. This new dataset comprises 318 configurations.
As for Classifier A, the dataset is split into 90% for training and 10% for validation (see Sec. II.2), and we optimize the model’s hyperparameters. The confusion matrices computed on the validation and complete set are shown in Fig. 7, while the classifier’s accuracy and $MCC$ are reported in Table 2. Although both $\alpha$ and $MCC$ are slightly lower than the values found for Classifier A, as expected given the increased complexity of the task, they are still large, above $90\%$.
@@ III.3 Classifier C
Finally, we create a classifier to distinguish between all four classes, employing the modified dataset as for Classifier B. We follow the same procedure, first creating the training and validation set, as explained in Sec. II.2, and then optimizing the model hyperparameters. From Fig. 8, showing the confusion matrices obtained on the validation set and the complete dataset, and Tab. 2, we see that the accuracy and $MCC$ are lower than the values found for the other two classifiers, as expected, but we still recover only a few misclassifications, with a validation accuracy $\sim 88\%$. In particular, we do not observe cases of “extreme” misclassification, i.e., in which a prompt collapse is classified as class IV or vice versa.
@@
IV Uncertainty in models’ predictions
In this section, we discuss how to quantify the uncertainty in our classifiers’ predictions.
For each input point $\mathbf{x}$, the classifier computes the probabilities for $\mathbf{x}$ to belong to each class $\omega_{k}$, $P(\omega_{k}|\mathbf{x})$, and its final prediction is that $\mathbf{x}$ belongs to the class with the largest probability. The likelihood of the preferred class provides the confidence, and can be used as a measure of the uncertainty of the prediction [112]. Figure 9 shows the distribution of the prediction confidence for Classifier A over the GW170817 samples. Regarding as an overall uncertainty estimate the mean of this distribution, and considering, for example, the results of Sec. III.1, we find that our model predicts GW170817 to have $60\%$ probability of formation of a NS remnant and $40\%$ prompt collapse with confidence $>99\%$. For the GW170817 samples obtained with the additional EoS and EM information (GW170817, GW170817+EoS, GW170817+EoS+KN, and GW170817+EoS+KN+GRB), we obtain a prediction confidence $\sim 99\%$ for Classifier A, $\sim 94\%$ for Classifier B, and $83-90\%$ for Classifier C. This estimate of uncertainty basically corresponds to quantifying “how sure” the classifier is about its prediction; therefore if two or more classes have very similar probabilities the uncertainty will be larger. This explains why Classifier B and Classifier C yield larger prediction uncertainties, especially in the latter case, where the classifier predicts basically the same probability for case II and case III for GW170817+EoS+KN+GRB (see Table 3).
@@ IV Uncertainty in models’ predictions
It is important to highlight that the largest source of uncertainty comes from the NR simulations employed to train the models. Different physics (e.g., neutrino treatments), settings, and resolutions can produce different $\tau_{\rm BH}$ even if the initial system configuration is the same. [^8] Examples of the effects of NR uncertainty can be found in Ref. [72], where the tables in Appendix C include the remnant classification obtained for different resolutions of the simulations. Additionally, Ref. [113], whose simulations are part of the CoRe database, shows, in Table II, the different collapse times obtained with different resolutions, and with or without simulating the effects of large magnetic fields with a subgrid model.
In the event of a detection, we need to know how much we can trust the prediction of these classifiers, especially if it informs EM follow-up. In this case, the input $\mathbf{x}$ consists of only one single point with the estimated parameter values provided by low-latency analyses, and not the full posterior obtained through the subsequent parameter estimation process. Therefore, apart from estimating the confidence of the prediction, as explained above, we can compute the probability of that point being misclassified to further decide whether to trust the prediction or not. We can calculate the uncertainty of the prediction for a point $\mathbf{x}$ based on the prediction entropy, and compare this value with a threshold found for the specific classifier employed: if the prediction uncertainty for $\mathbf{x}$ is smaller than the threshold, we trust the classification; if, on the other hand, the uncertainty is greater than the threshold, the probability of the prediction being a misclassification is non-negligible, and, therefore, more investigations about the event are necessary. In particular, for Classifiers A, B, and C, we find threshold values $T_{\rm\textsc{a}}\sim 0.3$, $T_{\rm\textsc{b}}\sim 0.06$, and $T_{\rm\textsc{c}}\sim 0.09$, respectively. The details about the uncertainty and the threshold calculations are explained in Appendix A.
@@ V Conclusions
We built a classifier to predict the outcome of a BNS merger based on parameters that can be measured from the inspiral GW signal, $M_{\rm tot}$, $q$, $\tilde{\Lambda}$, and $\chi_{\rm eff}$. We employed the Gradient Boosted Decision Tree algorithm in sklearn and a dataset comprising publicly available data from NR simulations. We constructed and trained three different models: Classifier A, to distinguish between prompt collapse to BH and formation of a NS remnant; Classifier B, to discriminate between prompt collapse, formation of a HMNS that collapses during the simulation, and formation of a NS remnant that does not collapse within the simulation time; Classifier C, which further discerns between short-lived ($2\,{\rm ms}<\tau_{\rm BH}<5\,{\rm ms}$) and long-lived ($\tau_{\rm BH}>5\,{\rm ms}$) HMNSs. All of them yielded very high values for both accuracy and the Matthews correlation coefficient (see Table 2), where the latter takes into account the distribution of correct classifications among the different classes. The lowest values obtained were $\alpha=88.2\%$ and $MCC=0.831$ for Classifier C, as expected given its more complex task.
@@
We investigated the outcome of GW170817 and GW190425 with the three classifiers, supplying as input the samples for $M_{\rm tot}$, $q$, $\tilde{\Lambda}$, and $\chi_{\rm eff}$ produced by parameter estimation analysis on the GW data. We concluded that GW170817 had a $\sim 60\%$ probability of forming a NS remnant, most likely a HMNS that collapsed in a few ms, while GW190425 collapsed promptly to a BH, with a $60-70\%$ probability. Moreover, we also applied our classifiers to the posterior samples obtained in Ref. [98], including in the inference information about the EoS and, in the case of GW170817, the EM counterpart. In this case, the probability for GW170817 to have formed a HMNS becomes larger than $99\%$, with roughly the same probability to be short- and long-lived, while GW190425 has more than $98\%$ probability to have undergone prompt collapse (see Table 3 for details).
@@
For all three classifiers, the most significant feature in determining the prediction is $\tilde{\Lambda}$, as expected since it is the parameter encoding the information about the EoS. Although the importance of $\chi_{\rm eff}$ is overall quite low, for Classifier A we find that, in general, binaries with $\chi_{\rm eff}<0$ ($\chi_{\rm eff}>0$) are more likely to produce a prompt collapse to BH (NS remnant) with respect to the irrotational case (see Fig. 6). However, only a few of the simulations included in our dataset corresponded to spinning configurations, and we did not consider misaligned spins, thus more data are needed to systematically investigate the spin influence.
@@
For Classifier A, which discriminates between prompt collapse and formation of a NS remnant, in the case of the GW170817 event we observed a clear threshold value for $\tilde{\Lambda}$ between the two scenarios (see Fig. 5). Specifically, for the samples median values of $M_{\rm tot}$, $q$, and $\chi_{\rm eff}$, we find that prompt collapse occurs for $\tilde{\Lambda}<\tilde{\Lambda}_{\rm th}\sim 310$.
@@
Finally, we discussed the uncertainty in the predictions of our models based on the probability assigned to each class, showing that our results on the real events data yielded a prediction confidence of roughly $90\%$ or higher.
@@
Knowing the remnant produced by a BNS merger could provide us with information about the EoS of NSs’ supranuclear-dense matter, the physical processes involved in the postmerger phase, and possible EM counterparts. Moreover, predicting whether a BNS system undergoes a prompt collapse to BH or produces a NS remnant is crucial to build postmerger models and to inform EM follow-up campaigns, especially for 3G detectors, for which a large number of BNS detections are expected. The classifiers developed in this work can be employed with any BNS detection, without the need for observing a postmerger signal. In fact, due to the high frequencies involved, the postmerger part of the GW signal is difficult to detect, and we expect to start observing it only with 3G detectors, but also in that case it could be weak [30, 33]. On the other hand, our models enable us to predict the merger outcome also for potential detections with Advanced Virgo and Advanced LIGO; moreover, with future-generation detectors, the more precise measurements of the binary parameters from the inspiral signal will allow us to get even more precise predictions, even without including additional information.